Angel A and angle B are complementary, that is their measurements add up to 90 degrees. Angle B measures 32 degrees more than angle A. What are the measurements of the two angles ?
Answer by TimothyLamb(4379)

b = a + 32
a + b = 90
---
a + b = 90
a + a + 32 = 90
2a = 58
---
a = 29
b = 61
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php